The half-life of 131I is 8.04 days. On a certain day, the activity of an iodine-131 sample is 6.36 mCi. What is its activity 37.8 days later?
Answer #1

A = A0*EXP(-lambda*t)

half life = 0.693/lambda

lambda = 0.693/8.04 = 0.0862

A = 6.36*10^-6*exp(-0.0862*37.8) = 0.245 mCi